show the complete solution




Student A investigated the rate of reaction between calcium carbonate (marble chips) and hydrochloric acid. The student collected 10.0 cm3 of gas produced after 15 seconds.Student B investigated the rate of reaction between magnesium and hydrochloric acid. The student collected 19.5 cm3 of gas produced after 10 seconds.Which student’s investigation had a faster rate of reaction?

1
Expert's answer
2022-04-20T14:54:04-0400

10.0 cm3 / 15 s = 0.667 cm3/s – first rate

19.5 cm3 / 10 s = 1.95 cm3 / s – second rate

1.95 / 0.667 = 2.9 times second rate is greater than the first